A sample of oxygen occupies a volume of 250. mL at a pressure of 704. torr. What volume will the gas occupy at a pressure of 800. torr if temerature is held constant
1
Expert's answer
2017-05-08T08:32:09-0400
P1×V1= P2×V2
V2 = P1×V1/P2
V2 = 704×250/800 = 220. ml

Answer:
220. ml.